C) What is the effective interest rate?
Answer #1

i = 20% / 4 = 5% per quarter

A. Payment each quarter = 1250 * (A/P,5%,5)

= 1250 * 0.05 * ((1 + 0.05)^5)/((1 + 0.05)^5-1)

= 1250 * 0.05 * (1.05^5)/(1.05^5-1)

= 1250 * 0.230975

= 288.72

B. Nominal interest rate = 20% compounded quarterly

C. Effective interest rate = (1+0.05)^4 -1

= 1.05^4 -1

= 1.2155 - 1

= 0.2155 ~ 21.55%

Pls comment if you think final solution is wrong or you require further explanation